# Defending Multimodal Backdoored Models by Repulsive Visual Prompt Tuning

### v2

@@ 4 Experiments
4.1 Experiment Settings
Models and benchmarks. Following the settings of [3], to create a more realistic backdoored CLIP for backdoor defense, we fine-tune OpenAI’s public checkpoint CLIP-400M [45] using 500K image-text pairs from the CC3M dataset [48], out of 1500 are poisoned with various types of backdoor attacks: BadNet [17], Blended [7], ISSBA [28], WaNet [41], TrojVQA [52] and BadCLIP [31]. We utilize ImageNet [10] to evaluate the performance of defense methods against all the aforementioned attacks, with the target class of the attacks set to banana. Additionally, we use Caltech101 [14] and OxfordPet [43] (a fine-grained dataset) to evaluate defense methods against BadNet, Blended, and WaNet. These attacks target “accordion” in Caltech101 and “samoyed” in OxfordPet, respectively. We compare RVPT with Linear Probe and CleanCLIP [3]. The implementation details of the backdoor attack methods, backdoor defense methods, and benchmarks will be illustrated in the supplementary material.
Implementation details. The implementation of our defense method is based on Pytorch [44]. Unless otherwise specified, we will use ViT/B32 [12] as CLIP’s visual encoder. We set the non-prompted depth $D=3$, learnable context length $b=50$, and balancing factor $\alpha=2$. Moreover, we set the proxy caption of the class the same as the simple prompt engineering in the original paper of CLIP [45]: (1) “a photo of <CLS>.” for ImageNet and Caltech101 (2) “a photo of <CLS>, a type of pet.” for OxfordPets. For all datasets, we randomly sample 16 images per class while setting the batch size and the epoch number to 32 and 50. The loss is optimized with SGD with an initial learning rate of 0.002 decayed by the cosine annealing rule. We conduct experiments on eight NVIDIA RTX 3090 GPUs and the computational expenses are shown in the supplementary material.
4.2 The Effectiveness of RVPT
@@
Firstly, we can see all attacks achieve a high ASR while maintaining decent CA with no defense, indicating their successful executions. The CA loss of the poisoned CLIP varies across downstream datasets because to obtain a backdoored CLIP, fine-tuning on the poisoned CC3M dataset [48] changes CLIP’s learned representations. These changes affect how CLIP performs on different datasets, accounting for varying levels of CA loss.
Table 3: Poisoned Accuracy (PA), defined as the model’s classification accuracy on poisoned images. Among these defense methods, RVPT achieves the highest PA, demonstrating its superior capability in preventing backdoored models from encoding the trigger, thereby maintaining accuracy on poisoned samples.
@@
Secondly, although CleanCLIP mitigates backdoor attacks in a zero-shot setting, it deteriorates CLIP’s recognition performance on downstream tasks and requires more computing resources (it needs full parameter tuning and more image-text pairs). Moreover, Linear Probe performs well against the first four attacks. However, it does not eliminate the trigger feature and thus performs the worst against BadCLIP, an attack specially designed for multimodal contrastive learning models. Since Linear Probe inevitably excludes CLIP’s text encoder, it cannot be used to defend against TrojVQA, which targets the multimodal task of visual question answering.
Finally, RVPT effectively defends against all types of attacks while maintaining clean accuracy. It demonstrates consistently strong defensive performance across all attacks and, notably, significantly reduces the ASR of BadCLIP [31], which is hard to defend by current defense methods. Moreover, as shown in Table 3, its classification accuracy on poisoned images is also much higher than other defense methods, highlighting its enhanced capability in maintaining correct predictions under backdoor attacks. Moreover, we observed that even a non-backdoored CLIP experiences a significant drop in PA, further supporting our claim of its low visual feature resistivity to input perturbation.
4.3 RVPT’s Defensive Ability Can Generalize
We evaluate RVPT’s generalization ability in defending backdoor attacks in three scenarios:
  • The target class of the attacks is absent from the tuning dataset but appears during testing.
  • The downstream dataset that will be attacked is different from the tuning dataset.
  • The downstream dataset that will be attacked shares different domains with the tuning dataset.
@@ 4.3 RVPT’s Defensive Ability Can Generalize
Method ImageNet-V2
BadNet Blended BadCLIP
No defense 86.55 (55.39) 99.04 (54.83) 99.89 (53.49)
@@
Method ImageNet-A
BadNet Blended BadCLIP
No defense 92.97 (31.47) 99.89 (31.22) 99.97 (30.80)
@@
Method ImageNet-R
BadNet Blended BadCLIP
No defense 66.63 (67.11) 97.94 (66.06) 99.69 (65.49)
@@
Method ImageNet-S
BadNet Blended BadCLIP
No defense 92.11 (41.73) 97.16 (41.86) 99.88 (40.19)
@@
Performance on the emerging target class.  We remove the samples belonging to the target class from the tuning dataset to evaluate RVPT’s performance on the emerging target class, which measures its defense capability for unseen target classes in the dataset, and the results are shown in Table 4. Moreover, we also tested BadCLIP (woTC) for the target class of banana and the ASR (CA) is 3.39 (61.89). Table 4 shows that for the backdoor attack of Blended and WaNet, RVPT performs even better in the absence of the target class, suggesting that RVPT has successfully filtered out the distinctive out-of-dataset pattern. Moreover, we can see that CA maintains other than OxfordPets, which decreases by around 2%. The CA drop is due to the dataset containing only 37 classes, so the absence of 3% of the data can considerably impact its accuracy. In summary, RVPT’s defensive capability effectively generalizes to attacks with emerging target classes that are not seen in the tuning dataset.
@@
Performance on cross-dataset transfer.  To evaluate RVPT’s defensive ability across datasets, we fine-tuned RVPT on ImageNet and tested it on Caltech101 and OxfordPets. The results are shown in Table 5. For all three attacks, although ASR increases slightly, it remains at a very low level. Additionally, RVPT’s accuracy decreases and occasionally falls below that of CleanCLIP. We conjecture the reason is that ImageNet does not share some of the in-dataset predictive features with Caltech101 or OxfordPets. In some cases, The non-essential feature unlearned in ImageNet may be the discriminative features in them and some of their non-essential features are not removed during tuning in ImageNet. Thus, there is a slight CA loss and ASR increase in these two datasets after RVPT. Nevertheless, the in-dataset predictive features of real-world images seem quite similar, ensuring RVPT’s cross-dataset defensive ability. Moreover, VPT’s tendency to overfit the base dataset is also one reason for the CA drop. Overall, RVPT’s performance on new datasets declines slightly but still defends better than CleanCLIP.
@@
Performance on cross-domain transfer.  To evaluate the cross-domain defensive ability, RVPT was trained on ImageNet and tested on ImageNet-V2 [47], ImageNet-A [19], ImageNet-R [18], and ImageNet-S [54]. The result is shown in Table 6. These three attacks can generalize well on the four benchmarks. We observe that RVPT will experience a loss of recognition and defense performance with domain shifts. In particular, it shows a significant CA drop on ImageNet-A, which reflects its effectiveness in focusing on in-dataset predictive features and ignoring other features, since ImageNet-A intentionally collects natural adversarial samples for ImageNet classifiers. Notably, unlike adversarial training [2], RVPT is not designed to defend against adversarial samples in ImageNet-A. The detailed comparison between RVPT and adversarial training is in the supplementary material.
@@ 4.4 Ablation Studies
Figure 3: ASR and CA are evaluated when the hyperparameters of shots, $\alpha$, context token length, depth, and margin are changed in RVPT. More ablation studies of visual encoder architecture and handcrafted text prompt can be referred to in the supplementary material.
Number of shots per class.  From Figure 3 (a), we surprisingly find that only 1 shot per class can make the ASR of Blended and BadNet attacks nearly 0 and significantly decrease the ASR of BadCLIP and increasing the number of shots further enhances the defense. For BadCLIP, 8 shots are needed to bring the ASR below 5%. Nevertheless, with 32 shots, the defense effectiveness stops improving.
Balancing factor $\alpha$.  We use $\alpha$ to balance the strengths of the CE loss and FR loss. From Figure 3 (b), we observe that both ASR and CA decrease as $\alpha$ increases. This observation indicates that if the balancing factor is set too high, it will interfere with the encoding of in-dataset predictive features (the adversarial environment is too harsh for CE loss optimization), resulting in performance degradation. However, a relatively small $\alpha$ is sufficient to achieve very low ASR while preserving strong CA for every evaluated dataset.
@@
Margin of the repelling strength.  To evaluate the potential risks of over-repelling, we set a margin $m$ of FR Loss $\mathcal{L}_{\mathrm{FRM}_{i}}$ for each sample $\boldsymbol{x}_{i}$:
$\mathcal{L}_{\mathrm{FRM}_{i}}=\mathrm{max}(\mathcal{L}_{\mathrm{FR}_{i}},m).$ (6)
A greater $m$ means a lower level of repulsion, and $m=1$ signifies no repulsion. The performance of RVPT of different $m$ is shown in Figure 3 (f). We observe that both ASR and CA will increase with larger $m$. Therefore, for the purpose of achieving superior defense performance, we adopt full repulsion that optimizes the mean cosine similarity to -1.
FR loss.  To assess the backdoor robustness introduced by FR loss in deep visual prompt tuning, we compare the ASR of RVPT and VPT against BadNet attack and Blended attack in different scenarios, as shown in Figure 4. The result indicates that while VPT alone reduces ASR to some extent, FR loss significantly enhances the defensive capability against backdoor attacks across various conditions.
@@
4.5 Qualitative Analysis
Figure 5: Attention map for (a) original image and (b) poisoned image. The second column shows the attention map of the backdoored visual encoder attacked by BadCLIP. The third and fourth columns show the attention map of RVPT’s patch embeddings and the prompt embeddings. RVPT effectively reduces the over-attention of both the non-essential features and the trigger pattern.
Attention map.  We plot the attention maps for the backdoored visual encoder and the defended visual encoder by RVPT in Figure 5 to observe the areas that the model focuses on. Firstly, for the clean image, RVPT enables the encoder to focus more on the useful features of the bird. In contrast, the backdoored encoder also attends to irrelevant features that humans cannot interpret. Secondly, for the poisoned image, although RVPT has noticed the trigger pattern, unlike the backdoored encoder, it will not draw all the attention to it.
t-SNE.  We draw the t-SNE plots [51] in Figure 6 to observe the visualizations of the image representations of different visual encoders. The plots show that image representations encoded by the backdoored CLIP and VPT-defended CLIP are clustered, indicating that these models focus solely on the trigger pattern, ignoring the information from the original image. On the contrary, RVPT not only disrupts this cluster but also restores the poisoned image representations, bringing them closer to their original counterparts. This observation suggests that after RVPT, the trigger pattern will have a minimal effect on the visual representations.
Figure 6: The t-SNE [51] plots for the representations of clean (dotted) and poisoned (star-shaped) images. The model is attacked by BadCLIP [31]. RVPT successfully breaks the clustering effect of the poisoned image features in the embedding space and ensures robust representations of attacked images, as observed in [3].
@@ 5 Conclusion
Supplementary Material
Organization of the Supplementary Material
  • Section A provides more details of the experimented perturbations and results of PR in CLIP backdoored by more attacks.
  • Section B provides more detail in the experiment’s settings of backdoor defense and attack methods.
  • Section C compares computational expenses between our method and CleanCLIP.
  • Section D contains more ablation studies, such as the visual encoder architecture, and the handcrafted prompt. It also contains the setting of experiments that ablates the FR loss.
  • Section E discusses the difference between adversarial training and the proposed RVPT.
A More Results of Perturbation Resistivity of More Backdoored CLIP
@@
The visualizations of these perturbations and trigger patterns can be found in Figure 7 and 10.
@@
The visual illustration for backdoored CLIP by BadNet and Blended are shown in Figure 8, 9 respectively. The results show that CLIP, even after VPT has generally much less PR value against various perturbations than traditionally supervised ViT, and RVPT can consistently increase its PR value across all scenarios.
@@ C Computational Expenses
D More Experiments for Ablation Study

### v3

@@ 4 Experiments
4.1 Experiment Settings
Models and benchmarks. Following the settings of bansal2023cleanclip , to create a more realistic backdoored CLIP for backdoor defense, we fine-tune OpenAI’s public checkpoint CLIP-400M radford2021clip using 500K image-text pairs from the CC3M dataset cc3m , out of 1500 are poisoned with various types of backdoor attacks: BadNet gu2019badnets , Blended chen2017blend , ISSBA issba , WaNet nguyen2021wanet , TrojVQA walmer2022dual and BadCLIP liang2024badclip . We utilize ImageNet deng2009imagenet to evaluate the performance of defense methods against all the aforementioned attacks, with the target class of the attacks set to banana. Additionally, we use Caltech101 fei2004learning and OxfordPet parkhi2012cats (a fine-grained dataset) to evaluate defense methods against BadNet, Blended, and WaNet. These attacks target “accordion” in Caltech101 and “samoyed” in OxfordPet, respectively. We compare RVPT with Linear Probe and CleanCLIP bansal2023cleanclip . The implementation details of the backdoor attack methods, backdoor defense methods, and benchmarks will be illustrated in Appendix B.
Implementation details. Unless otherwise specified, we will use ViT/B32 dosovitskiy2020vit as CLIP’s visual encoder. The prompt is initialized from a zero mean Gaussian distribution. We set the non-prompted depth $D=3$, learnable context length $b=50$, and balancing factor $\alpha=2$. Moreover, the proxy caption of the class the same as the simple prompt engineering in the original paper of CLIP radford2021clip : (1) “a photo of <CLS>.” for ImageNet and Caltech101 (2) “a photo of <CLS>, a type of pet.” for OxfordPets. For all datasets, we randomly sample 16 images per class while setting the batch size and the epoch number to 32 and 50. The loss is optimized with SGD with an initial learning rate of 0.002 decayed by the cosine annealing rule. We conduct experiments on eight NVIDIA RTX 3090 GPUs and the computational expenses are shown in Appendix C.
4.2 The Effectiveness of RVPT
@@
Firstly, we can see all attacks achieve a high ASR while maintaining decent CA with no defense, indicating their successful executions. The CA loss of the poisoned CLIP varies across downstream datasets because to obtain a backdoored CLIP, fine-tuning on the poisoned CC3M dataset cc3m changes CLIP’s learned representations. These changes affect how CLIP performs on different datasets, accounting for varying levels of CA loss.
Secondly, although CleanCLIP mitigates backdoor attacks in a zero-shot setting, it deteriorates CLIP’s recognition performance and requires more computing resources. Moreover, Linear Probe performs well against the first four attacks. However, it does not eliminate the trigger feature and thus performs the worst against BadCLIP, an attack specially designed for multimodal contrastive learning models. Since Linear Probe inevitably excludes CLIP’s text encoder, it cannot be used to defend against TrojVQA, which targets the multimodal task of visual question answering.
[…]
@@
Finally, RVPT effectively defends against all types of attacks while maintaining clean accuracy. It demonstrates consistently strong defensive performance across all attacks and, notably, significantly reduces the ASR of BadCLIP liang2024badclip , which is hard to defend by current defense methods. Moreover, as shown in Table 3, RVPT’s classification accuracy on poisoned images is also much higher than other defense methods, highlighting its enhanced capability in maintaining correct predictions under backdoor attacks. Moreover, we observed that even a non-backdoored CLIP experiences a significant drop in PA, further supporting our claim of its low visual feature resistivity to input perturbation.
@@
      Dataset       Method       BadNet       Blended       BadCLIP
      ImageNet-V2       No defense       86.55 (55.39)       99.04 (54.83)       99.89 (53.49)
@@
      ImageNet-A       No defense       92.97 (31.47)       99.89 (31.22)       99.97 (30.80)
@@
      ImageNet-R       No defense       66.63 (67.11)       97.94 (66.06)       99.69 (65.49)
@@
      ImageNet-S       No defense       92.11 (41.73)       97.16 (41.86)       99.88 (40.19)
@@
4.3 The Generalization of RVPT’s Defensive Ability
We evaluate RVPT’s generalization ability in defending backdoor attacks in three scenarios: (1) The target class of the attacks is absent from the tuning dataset but appears during testing. (2) The downstream dataset that will be attacked is different from the tuning dataset. (3) The downstream dataset that will be attacked shares different domains with the tuning dataset.
@@ 4.3 The Generalization of RVPT’s Defensive Ability
Performance on the emerging target class. We remove the samples belonging to the target class from the tuning dataset to evaluate RVPT’s performance on the emerging target class, which measures its defense capability for unseen target classes in the dataset, and the results are shown in Table 4. Moreover, we also tested BadCLIP (woTC) for the target class of banana and the ASR (CA) is 3.39 (61.89). Table 4 shows that for the backdoor attack of Blended and WaNet, RVPT performs even better in the absence of the target class, suggesting that RVPT has successfully filtered out the distinctive out-of-dataset pattern. Moreover, we can see that CA maintains except OxfordPets, which decreases by around 2%. The CA drop is due to this dataset containing only 37 classes, so the absence of 3% of the data can impact its accuracy. In summary, RVPT’s defensive capability effectively generalizes to attacks with emerging target classes that are not seen in the tuning dataset.
@@
Performance on cross-dataset transfer. To evaluate RVPT’s defensive ability across datasets, we tuned CLIP with RVPT on ImageNet and tested it on Caltech101 and OxfordPets. The results are shown in Table 5. For all three attacks, although ASR increases slightly, it remains at a very low level. Additionally, RVPT’s accuracy decreases and occasionally falls below that of CleanCLIP. We conjecture the reason is that ImageNet does not share some of the in-dataset predictive features with Caltech101 or OxfordPets. In some cases, The non-essential feature unlearned in ImageNet may be the discriminative features in them and some of their non-essential features are not removed during tuning in ImageNet. Thus, there is a slight CA loss and ASR increase after RVPT. Nevertheless, the in-dataset predictive features of real-world images seem quite similar, ensuring RVPT’s cross-dataset defensive ability. Moreover, VPT’s tendency to overfit the base dataset also contributes to the CA drop. Overall, RVPT’s performance on new datasets declines slightly but still outperforms CleanCLIP.
@@
Performance on cross-domain transfer. To evaluate the cross-domain defensive ability, RVPT was trained on ImageNet and tested on ImageNet-V2 imegenetv2 , ImageNet-A imageneta , ImageNet-R imagenetr , and ImageNet-S imagenets . The result is shown in Table 6. RVPT can generalize its defensive ability across different domains. We also observe that RVPT will experience a loss of recognition and defense performance with domain shifts. In particular, it shows a significant CA drop on ImageNet-A, which reflects its effectiveness in focusing on in-dataset predictive features and ignoring other features, since ImageNet-A intentionally collects natural adversarial samples for ImageNet classifiers. Notably, unlike adversarial training bai2021recent , RVPT is not designed to defend against adversarial samples in ImageNet-A. The detailed comparison between RVPT and adversarial training is in Appendix E.
@@ 4.4 Ablation Studies
Figure 3: ASR and CA are evaluated when the hyperparameters of shots, $\alpha$, context token length, depth, and margin are changed in RVPT. In panel (f), the dashed lines represent the situation when the total FR loss is ablated. More ablation studies of visual encoder architecture and handcrafted text prompt can be referred to in Appendix D.
Number of shots per class. From Figure 3(a), we surprisingly find that only 1 shot can make the ASR of Blended and BadNet attacks nearly 0 and significantly decrease the ASR of BadCLIP and increasing the number of shots further enhances the defense. For BadCLIP, 8 shots are needed to bring the ASR below 5%. Nevertheless, with 32 shots, the defense effectiveness stops improving.
Balancing factor $\alpha$. We use $\alpha$ to balance the strengths of the CE loss and FR loss. From Figure 3(b), we observe that both ASR and CA decrease as $\alpha$ increases. This observation indicates that if the balancing factor is set too high, it will interfere with the encoding of in-dataset predictive features (the adversarial environment is too harsh for CE loss optimization), resulting in performance degradation. However, a relatively small $\alpha$ is sufficient to achieve very low ASR for every evaluated dataset.
@@
Margin of the repelling strength. To evaluate the potential risks of over-repelling, we set a margin $m$ of FR Loss $\mathcal{L}_{\mathrm{FRM}_{i}}$ for each sample $\boldsymbol{x}_{i}$: $\mathcal{L}_{\mathrm{FRM}_{i}}=\mathrm{max}(\mathcal{L}_{\mathrm{FR}_{i}},m)$. A greater $m$ means a lower level of repulsion, and $m=1$ signifies no repulsion (which is VPT). The performance of RVPT of different $m$ is shown in Figure 3(f). We observe that both ASR and CA will increase with larger $m$. Therefore, for the purpose of achieving superior defense performance, we adopt full repulsion that optimizes the mean cosine similarity to -1.
4.5 Further Explorations
Figure 4: Last-layer attention map for (a) original (b) poisoned image in backdoored model (attacked by BadCLIP) and RVPT.
Figure 5: The t-SNE JMLR:v9:vandermaaten08a plots for the representations of clean (dotted) and poisoned (star-shaped) images (attacked by BadCLIP).
Attention map and t-SNE. We analyze the attention maps and t-SNE JMLR:v9:vandermaaten08a plots of different models. In Figure 5, RVPT enables the encoder to focus on relevant bird features for clean images, while the backdoored model attends to irrelevant ones. For poisoned images, RVPT notices the trigger but avoids overemphasizing it, unlike the backdoored encoder. Figure 5 shows that backdoored CLIP and VPT models cluster around the trigger pattern, ignoring original image content bansal2023cleanclip . RVPT, however, disrupts this cluster and restores poisoned image representations, reducing the influence of the trigger.
RVPT endows CLIP with backdoor robustness in various downstream tasks. To further evaluate the impact of backdoored CLIP in real-world settings, we assess its performance on two downstream tasks: image captioning and text-to-image retrieval in Table 7. For image captioning, we use MSCOCO lin2014microsoft and LLaVA-1.5 (13B) liu2023llava equipped with backdoored CLIP ViT/L14. A successful attack is defined as generating a caption that includes the target class. For text-to-image retrieval, we use an enriched ImageNet-1K deng2009imagenet ; huggingface2025vl with backdoored CLIP (ViT/B32). We insert one poisoned image into the candidate set and if it appears in the top-10 matches for clean captions of the target class, we count it as a successful attack. Table 7 shows that backdooring CLIP during pre-training can transfer the backdoor behavior to the evaluated downstream tasks. Notably, our RVPT effectively mitigates backdoor effects while preserving clean performance, outperforming baselines bansal2023cleanclip ; jia2022visual .
@@ 4.5 Further Explorations
Image Captioning (IC) Text-to-Image Retrieval (TR)
[…]
Clean model - 7.65 - 95.83 69.34
[…]
CleanCLIP 23.03 4.32 13.33 25.54 68.95
VPT 81.21 4.64 22.67 26.80 71.88
RVPT 0 4.51 0 82.12 69.89
@@ 5 Conclusion
Appendix
We summarize the Appendix as follows:
  • Appendix A provides more details of the experimented perturbations and results of PR in CLIP backdoored by more attacks.
  • Appendix B provides more detail in the experiment’s settings of backdoor defense and attack methods.
    • Section B.1 provide the detailed settings of backdoor attacks.
    • Section B.2 provide the detailed settings of backdoor defense baselines.
    • Section B.3 provide the detailed settings of adopted datasets.
  • Appendix C compares computational expenses between our method and CleanCLIP.
  • Appendix D contains more ablation studies, such as the visual encoder architecture, and the handcrafted prompt. It also contains the setting of experiments that ablates the FR loss.
    • Section D.1 provide the ablation study for handcrafted prompt.
    • Section D.2 provide the ablation study for visual encoder architecture.
  • Appendix F discusses both the positive and negative of potential impact of our RVPT.
  • Appendix G discusses limitations of the proposed RVPT.
Appendix A More Results of Perturbation Resistivity of More Backdoored CLIP
@@
The visualizations of these perturbations and trigger patterns can be found in Figure 6 and 9.
@@
  • ViT/B32 supervised exclusively on ImageNet.
The visual illustration for backdoored CLIP by BadNet and Blended are shown in Figure 7, 8 respectively. The results show that CLIP, even after VPT has generally much less PR value against various perturbations than traditionally supervised ViT, and RVPT can consistently increase its PR value across all scenarios.
@@ Appendix C Computational Expenses
Appendix D More Experiments for Ablation Study